For what values of c will the system have:
infinitely many solution
1) 2x - 6y = c
6x - 18y = 30
2)no solution
x + 2y = 6
cx - 4y = 8


How do you do these questions? Can you explain how?

Answer by vleith(2517)

infinitely many solution
1) 2x - 6y = c
6x - 18y = 30
infinitely many solution means the two 'lines' are really the same line
the second line above has two terms that are exactly 3 times the terms of the fisrt line (6x = 3*2x and -18y = 3 *(-6y)) so if c is 3*30, then the two lines are really the same line.
2)no solution
x + 2y = 6
cx - 4y = 8
No solution occurs when the 'answer' is something that can't happen
So let's double the first line to get
2x + 4y = 12
Now add the second line to it
2x + 4y = 12
cx -4y = 8
----------
(2+c)x = 20
any c that that makes (2+x) a non-zero number will result in an x that 'works'. However, if c = -2, then we get
(0*x) = 20
0 times anything is 0, so that yields
0 = 20
0=20 is impossible, so is c = -2 then ther eis no solution.,
make sense?